OpenStudy (anonymous):

how do you solve 11=43.6e^.1947t

OpenStudy (anonymous):

Ln(11/43.6)=0.194t t=(Ln(11/43.6))/(0.194)

OpenStudy (campbell_st):

well 1st \[11/43.6 = e^{0.1947t}\] next take the base e log of each side so ln (11/43.6) = 0.1947t then t = ln(11/43/6)/0.1947